OpenStudy (lgbasallote):

how to turn \[\frac{1}{s^3(s^2+1)}\] into partial fractions? im having difficulties :/

OpenStudy (experimentx):



OpenStudy (turingtest):

\[\frac1{s^3(s^2+1)}=\frac As+\frac B{s^2}+\frac C{s^3}+\frac{Dx+E}{s^2+1}\]so far okay @lgbasallote ?

OpenStudy (lgbasallote):

wait...in partial fractions dont i do \[\frac As + \frac{B}{s^2} + \frac{C}{s^3}?\]

OpenStudy (lgbasallote):

yes @TuringTest that's what i did

OpenStudy (turingtest):

so you agree with my line lgb\[\frac1{s^3(s^2+1)}=\frac As+\frac B{s^2}+\frac C{s^3}+\frac{Dx+E}{s^2+1}\]now multiply both sides by \(s^3(s^2+1)\) and see what you get

OpenStudy (lgbasallote):

i got As^2 (s^2 + 1)( + Bs(s^2 + 1) + C(s^2 + 1) + Ds (s^3) + Es^3

OpenStudy (lgbasallote):

=1

OpenStudy (turingtest):

yep, now collect like terms

OpenStudy (lgbasallote):

by root sub i got when s = 0; C = 1 after that i got nothing

OpenStudy (lgbasallote):

i tried gaussian-ing...but no juice

OpenStudy (turingtest):

it's one or the other, and since you have most variables attatched to an s you really do have to do the gaussian way I think

OpenStudy (turingtest):

\[As^4+ As^2+ Bs^3 + Bs+ Cs^2 + C + Ds^4 + Es^3=1\]let me know if you see a mistake...

OpenStudy (lgbasallote):

yup same as mine...

OpenStudy (lgbasallote):

i now got D = 0

OpenStudy (turingtest):

\[(A+D)s^4+ (B+E)s^3+ (A+C)s^2 + Bs + C =1\]still the same?

OpenStudy (lgbasallote):

wait...how did you get ss^4 in a o.O

OpenStudy (lgbasallote):

uhh darn messy solution..i see it now

OpenStudy (lgbasallote):

okay now i have A = -1 and D = 1 im stuck with B and E

OpenStudy (turingtest):

let me try to reason it out C=1 A+C=0 -> A=-1 A+D=0 -> D=1 B=0 (since the coefficient of s is 0) so B+E=0 -> E=0

OpenStudy (lgbasallote):

.....oh

OpenStudy (lgbasallote):

of course...

OpenStudy (turingtest):

you didn't realize that you knew B I've made that mistake wayyyyyy to many times to do it again ! (that's what I tell myself every time at least)